During startup of a power plant, a turbine accelerates from rest at 0.58 rad/s2 .

1-How long does it take to reach its 2600 rpm operating speed? Express your answer using two significant figures

2-How many revolutions does it make during this time?Express your answer using two significant figures.
Answer #1

1.

wo = initial angular speed = 0 rad/s

wf = final angular speed = 2600 rpm = 2600 (2pi/60) = 272.13 rad/s

t = time taken = ?

\alpha = angular acceleration = 0.58

using the equation

wf = wo + \alpha t

272.13 = 0 + (0.58) t

t = 469.2 sec

2)

\theta = angular displacement = wo t + (0.5) \alpha t2 = (0) (469.2) + (0.5) (0.58) (469.2)2 = 6.4 x 104

n = number of revolution = \theta/2pi = (6.4 x 104)/(2 x 3.14) = 1.02 x 104 revolution